Buffer solutions:


1. Calculate the pH of a carbonate buffer system that contains 0.0065 M HCO3- and 0.0084 M CO32-. The Ka of HCO3- is 4.7 X 10-11.



1
Expert's answer
2021-04-15T03:22:17-0400

PH=pKa+log(salt/acid)(pKa=-log(ka)]

pH= 10.32+log(0.0084/0.0065)

pH=10.32+0.11

pH= 10.43